OpenStudy (vuriffy):

\[\sin (2x) = -\sqrt{(2)}/(2) \]

OpenStudy (vuriffy):

Solve each equation for 0 <= x <2pi

OpenStudy (sshayer):

\[\sin 2x=-\frac{ \sqrt{2} }{ 2 }=-\frac{ 1 }{ \sqrt{2} }=-\sin \frac{ \pi }{ 6 }=\sin \left( \pi+\frac{ \pi }{ 6 } \right),\sin (2 \pi- \frac{ \pi }{6 } ),\] \[\sin \left( 2 \pi+\pi+\frac{ \pi }{ 6 } \right),\sin \left( 2 \pi+2 \pi-\frac{ \pi }{ 6 } \right)\] 2x=? x=?

OpenStudy (vuriffy):

I don't really understand this at all. I do under the first 3 steps though.

OpenStudy (vuriffy):

understand*

OpenStudy (vuriffy):

Wifi dropped, sorry..

OpenStudy (vuriffy):

Is there only two solutions?

OpenStudy (vuriffy):

I know you can make a graph for these, but I have never made a sin(2x) graph

OpenStudy (sshayer):

sorry i wrongly typed

OpenStudy (sshayer):

\[2x=\frac{ 5 \pi }{ 4 },\frac{ 7 \pi }{ 4 },\frac{ 13 \pi }{ 4 },\frac{ 15 \pi }{ 4 }\] x=?

OpenStudy (vuriffy):

Am I supposed to divide by 2?

OpenStudy (sshayer):

write \[\frac{ \pi }{ 4 }~\in~place~of~\frac{ \pi }{ 6 } \]

OpenStudy (sshayer):

yes divide by 2

OpenStudy (vuriffy):

I don't think I saw one before where it went from 5 (+2), 7 (+6), 13(+2), Isn't it a constant?

OpenStudy (sshayer):

if it is 2x ,then add 2n pi,where n=0,1 if it is 3x,then add 2 n pi,where n=0,1,2 .....

OpenStudy (vuriffy):

I'm new to this subject, so I haven't fully grasped it yet.

OpenStudy (sshayer):

\[\sin 2x=\sin \left( 2 n \pi+\frac{ 5 \pi }{ 4 } \right),\sin \left( 2n \pi+\frac{ 7 \pi }{ 4 } \right)\]

OpenStudy (vuriffy):

one answer is pi/4?

OpenStudy (sshayer):

as sin 2x is negative so it lies in 3 rd and 4th quadrant.

OpenStudy (vuriffy):

All the solutions are in the 3rd and 4th?

OpenStudy (sshayer):

\[\frac{ 5 \pi }{ 8 },\frac{ 7 \pi }{ 8 },\frac{ 13 \pi }{ 8 },\frac{ 15 \pi }{ 8 }\]

OpenStudy (vuriffy):

Does that mean there is 8 solutions?

OpenStudy (sshayer):

no ,we have to take between 0 and 2 pi . they are only four.

OpenStudy (vuriffy):

Oh, I see now.

OpenStudy (vuriffy):

Do you know what the graph looks like for this kind of equation?

OpenStudy (sshayer):

remember 2x lies in 3rd and 4th quadrant.

OpenStudy (vuriffy):

So the negative part of the quadrant, yes?

OpenStudy (sshayer):

sorry i have to go to bed now and leaving.